OpenStudy (anonymous):

if i drop two objects from the same height will they hit the ground at the same time if they are different wieight?

OpenStudy (anonymous):

yes both object hit the ground at same time (We need to ignore air resistance for lighter objects).

OpenStudy (anonymous):

gravitational pull objects at the same rate of 9.8 meters per second

OpenStudy (anonymous):

The velocity when an object falls is (2gh)^0.5 Where g is gravity and h is height We see that velocity of fall is independent of mass of the body

OpenStudy (harsha19111999):

Actually, the heavier body reaches ground first. For example, if you take a piece of paper and a ball. The ball reaches the ground first. But, if you do not consider air drag(resistance), definitely they both will reach the ground at the same time

OpenStudy (anonymous):

When an object is on the height ''h'' and it is dropped down because of the gravity without the effect of any other power from that experiment's environment, then when this object hits the ground then that has a velocity. In accordance with the law of conservation of energy, the total energy of that object on the height ''h'' will be equal with the total energy at the point ''A'' in which the object hits the ground. Therefore on the height ''h'' the total energy is Ek (h) (kinetic energy) + Ed (h) (dynamic energy). And at the point ''A'' the total energy is Ek (A) (kinetic energy) + Ed (A) (dynamic energy). According to the law of conservation of energy is valid that: Ek (h) + Ed (h) = Ek (A) + Ed (A) (1) Ek (h) = 0 (2) Ed (h) = m*g*h (3) Ek (A) = \[\frac{ 1 }{ 2 } * m*u ^{2}\] (4) Ed (A) = 0 (5) Then solving the system of the above 5 equations as to "u" which is the velocity of the object when that hits the ground at the point ''A'', it is found out that: \[u = \sqrt{2*g*h}\] Where ''g'' is the acceleration of the gravity. Then it is observed that any object has velocity ''u'' into any environment in which the one and only power on it is that of the power of the gravity and the velocity ''u'' is independent of the mass that means the weight does not take place in that phenomena and it does not make sense to compare the weights of the objects in that case.